OpenStudy (anonymous):

The first aircraft has 87 more seats than the second aircraft. The third aircraft has 25 fewer seats than the second aircraft. If their total number of seats is 386, find the number of seats for each aircraft

OpenStudy (texaschic101):

f = 1st aircraft s = 2nd aircraft t = 3rd aircraft total number of sears is 386..... f + s + t = 386 1st aircraft has 87 more seats then second.. f = s + 87 3rd aircraft had 25 fewer seats then 2nd... t = s - 25 now we are going to sub s - 25 in for t, and s + 87 in for f, back into the first equation f + s + t = 386 (s + 87) + s + (s - 25) = 386 --- combine like terms 3s + 62 = 386 3s = 386 - 62 3s = 324 s = 324/3 s = 108 so if the 2nd aircraft has 108, then... f = s + 87 = 108 + 87 = 195 (seats in 1st aircraft) t = s - 25 = 108 - 25 = 83 (seats for 3rd aircraft) check... f + s + t = 386 195 + 108 + 83 = 386 386 = 386 (correct)

OpenStudy (anonymous):

Thank you so much, man.